# CTC-DRO: Robust Optimization for Reducing Language Disparities in Speech Recognition

### v2

@@ 2 Cantonese yue CV, Fleurs
Table 6: Dataset statistics for the training set of each of the language subsets used in our experiments, in the balanced data setting. ISO codes are used for the languages, duration is presented in seconds, and transcript length is in number of characters. Averages and standard deviations are reported.
Set # ISO Number of Duration Transcript
Data Points Length
1 ces 908 $4.0\pm 1.7$ $23.8\pm 22.1$
cmn 322 $10.4\pm 3.5$ $36.8\pm 13.9$
nan 1406 $2.6\pm 0.7$ $3.4\pm 1.9$
pol 482 $7.5\pm 3.0$ $104.6\pm 46.3$
ron 274 $12.6\pm 3.1$ $136.1\pm 45.1$
spa 445 $8.1\pm 2.2$ $91.1\pm 26.4$
2 eng 647 $4.7\pm 1.5$ $63.7\pm 25.4$
fas 693 $5.2\pm 1.7$ $34.4\pm 18.2$
hrv 291 $11.7\pm 3.3$ $116.3\pm 35.7$
ita 326 $10.7\pm 3.2$ $140.4\pm 42.3$
slk 330 $10.6\pm 3.3$ $116.2\pm 38.6$
yue 243 $12.2\pm 3.7$ $31.7\pm 10$2
3 khm 206 $13.7\pm 3.4$ $122.5\pm 36.5$
kmr 723 $5.0\pm 1.6$ $30.8\pm 15.0$
kor 269 $12.5\pm 3.0$ $45.8\pm 14.1$
nbl 744 $4.8\pm 1.9$ $31.3\pm 10.0$
nno 709 $4.5\pm 1.2$ $41.2\pm 17.3$
tat 835 $4.3\pm 1.8$ $33.2\pm 20.8$
4 mrj 707 $5.1\pm 2.0$ $40.8\pm 22.8$
slv 918 $3.9\pm 1.1$ $30.2\pm 12.3$
snd 263 $12.0\pm 3.6$ $105.4\pm 31.2$
sot 655 $5.5\pm 2.0$ $51.0\pm 23.6$
spa 550 $6.6\pm 3.4$ $87.2\pm 50.2$
urd 299 $11.3\pm 3.4$ $119.9\pm 37.1$
5 deu 745 $4.8\pm 1.6$ $43.3\pm 16.1$
eng 712 $5.0\pm 1.5$ $47.7\pm 17.4$
heb 345 $10.2\pm 3.3$ $91.9\pm 29.8$
jpn 290 $11.5\pm 3.1$ $50.0\pm 15.8$
rus 318 $10.8\pm 3.4$ $125.6\pm 42.2$
spa 311 $11.1\pm 3.4$ $144.6\pm 50.0$
Appendix B Results
For each language set, we present the language-specific results of our experiments using balanced training data in Table 7. Table 8 shows the language-specific results for the first two sets based on experiments using all available training data in ML-SUPERB 2.0.
@@ Appendix B Results
Table 7: Results of the baseline models, group DRO models, and CTC-DRO models on the different language sets, where languages are indicated by their ISO code. We show the CER on the individual languages, CER averaged across languages (Avg CER), and LID accuracy (LID) for fine-tuned MMS and XLS-R models. Best LID and CER results are highlighted, and the CERs for the worst-performing languages are underlined.
@@
1 MMS Baseline 8.4 52.4 60.8 3.6 13.3 1.8 23.4 97.4
group DRO 20.6 48.6 86.6 4.3 16.7 6.2 30.5 78.7
CTC-DRO 10.5 46.1 56.8 3.7 17.9 2.3 22.9 95.8
XLS-R Baseline 7.3 64.9 60.8 3.1 13.4 1.8 25.2 92.6
group DRO 27.4 48.9 78.4 3.7 14.9 6.6 30.0 87.8
CTC-DRO 7.8 50.7 57.6 3.0 14.2 1.8 22.5 89.5
@@
2 MMS Baseline 0.2 21.8 9.0 5.9 8.2 49.4 15.8 98.4
group DRO 11.8 29.7 10.8 6.2 10.2 55.5 20.7 98.2
CTC-DRO 0.5 22.1 8.8 5.5 8.6 44.4 15.0 96.2
XLS-R Baseline 0.1 20.6 10.9 4.6 8.9 68.8 19.0 94.2
group DRO 12.7 28.5 14.4 5.1 10.2 58.8 21.6 87.0
CTC-DRO 0.5 21.5 12.6 5.2 10.0 45.0 15.8 89.3
@@
3 MMS Baseline 31.3 12.2 34.2 7.4 2.5 9.0 16.1 98.5
group DRO 33.2 19.1 34.0 22.4 9.8 13.5 22.0 98.7
CTC-DRO 31.3 12.0 27.6 8.1 2.3 10.2 15.3 98.7
XLS-R Baseline 33.2 13.3 32.3 8.7 3.7 11.0 17.0 99.2
group DRO 38.0 23.9 35.5 26.6 11.9 14.9 25.1 97.2
CTC-DRO 32.2 14.8 31.9 10.1 5.0 12.0 17.7 97.9
@@
4 MMS Baseline 14.8 6.9 24.0 14.4 5.9 20.1 14.4 87.9
group DRO 13.1 14.4 19.0 17.1 3.8 21.8 14.9 91.9
CTC-DRO 17.7 8.1 17.5 11.4 4.4 18.4 12.9 87.3
XLS-R Baseline 14.0 4.8 23.3 11.6 4.2 29.7 14.6 88.4
group DRO 19.5 25.6 18.5 23.0 3.9 21.1 18.6 83.5
CTC-DRO 11.9 6.7 21.0 13.8 4.8 24.2 13.7 88.9
@@
5 MMS Baseline 5.4 11.1 30.2 90.0 12.0 7.2 26.0 96.3
group DRO 27.6 27.0 32.6 62.2 17.6 8.4 29.2 67.0
CTC-DRO 10.9 15.4 39.2 57.5 13.2 9.3 24.3 90.5
XLS-R Baseline 4.8 9.2 33.2 114.8 10.5 7.1 29.9 89.0
group DRO 29.1 26.8 46.1 92.9 16.5 9.3 36.8 57.7
CTC-DRO 5.7 9.6 38.6 71.5 10.1 7.3 23.8 91.0
Table 8: Results of the baseline models, group DRO models, and CTC-DRO models on the first two language sets using additional amounts of training data per language, where languages are indicated by their ISO code. We show the CER on the individual languages, CER averaged across languages (Avg CER), and LID accuracy (LID) for fine-tuned MMS and XLS-R models. Best LID and CER results are highlighted, and the CERs for the worst-performing languages are underlined.
@@
1 MMS Baseline 9.1 58.9 67.5 6.0 7.1 5.0 25.6 98.1
group DRO 13.8 92.1 96.3 6.7 11.9 5.8 37.8 83.9
CTC-DRO 8.7 45.9 62.8 6.2 7.5 5.3 22.8 98.5
XLS-R Baseline 13.0 92.1 78.3 9.8 12.0 8.5 35.6 96.4
group DRO 18.9 86.4 90.8 5.7 21.6 5.0 38.1 72.3
CTC-DRO 12.9 52.5 67.5 9.0 11.9 7.8 26.9 97.1
@@
2 MMS Baseline 9.6 16.9 8.5 6.8 8.0 66.9 19.5 99.0
group DRO 10.1 70.0 24.3 7.9 14.8 105.4 38.8 81.0
CTC-DRO 9.7 18.1 8.3 6.6 7.3 48.1 16.4 99.1
XLS-R Baseline 11.9 32.2 9.6 8.1 9.2 97.2 28.0 98.2
group DRO 8.8 88.2 33.9 6.7 23.3 102.9 44.0 80.8
CTC-DRO 11.6 23.2 9.3 8.2 8.9 51.4 18.8 98.6
Appendix C Ablation Study

### v3

@@ 2 Cantonese yue CV, Fleurs
Appendix E Evaluation Metric Details
In Section 4, we discuss the evaluation metrics used. Here, we provide more details about the computation of the CER. The CER can be computed by comparing the system generated and reference transcripts using the formula:
$\text{CER}=\frac{I+S+D}{N}\times 100,$ (18)
where $I$ is the number of insertions, $S$ the number of substitutions, and $D$ the number of deletions in a minimum edit distance alignment between the reference and system output, and $N$ is the number of characters in the reference transcript. The WER is computed identically, but operates at the word level rather than the character level (see WER results in Appendix F.2).
Appendix F Results
In Section F.1, we present the language-specific results on the development set, showing the effect of our tested hyperparameters. In addition, we show the language-specific test results in Section F.2. In this section, we include a WER analysis for set 4 for completeness. This set was chosen, as it contains languages with clear word boundaries. Additionally, we present the language-specific results of our ablation study and an analysis of the batch duration hyperparameter in Section F.3. We present multi-seed experiments on sets 1 and 3 in Section F.4. Finally, we address the effect on the best-performing language in Section F.5 and plot the group weights for additional language sets and models in Section F.6.
F.1 Language-Specific Development Results
To show the effect of our tested hyperparameters on the performance of the CTC-DRO models, we present language-specific results on the development set. In Table 7, we show the development results for tested values of $\eta_{q}\in\{10^{-3},10^{-4}\}$ and $\alpha\in\{0.1,0.5,1\}$ in the balanced data setup. The results for models trained with additional training data are shown in Table 8. For each language set, the model with the best-performing hyperparameter setting is evaluated on the test data. All results are obtained using a learning rate of $10^{-4}$.
@@ F.1 Language-Specific Development Results
Table 7: Results of the CTC-DRO models on the development set for the different language sets, where languages are indicated by their ISO code. We show the CER on the individual languages and CER averaged across languages (Avg) for fine-tuned MMS and XLS-R models. We highlight the best hyperparameter setting per set.
Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ ces cmn nan pol ron spa Avg Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ eng fas hrv ita slk yue Avg
# ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) # ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
1 MMS $10^{-3}$ 0.1 11.6 45.5 58.7 4.2 16.0 2.6 23.1 2 MMS $10^{-3}$ 0.1 0.4 28.8 8.1 5.0 8.7 48.5 16.6
$10^{-3}$ 0.5 12.4 44.4 56.4 4.4 16.9 3.0 22.9 $\bm{10^{-3}}$ 0.5 0.6 30.0 8.0 5.2 7.8 44.8 16.0
$10^{-3}$ 1.0 12.2 45.8 56.1 4.5 16.6 2.7 23.0 $10^{-3}$ 1.0 1.1 28.0 7.9 5.0 8.9 45.2 16.0
$10^{-4}$ 0.1 9.8 49.1 62.2 4.2 16.4 2.5 24.0 $10^{-4}$ 0.1 0.4 27.3 7.1 5.0 7.4 46.8 15.6
$10^{-4}$ 0.5 13.0 47.7 61.8 4.3 17.4 2.8 24.5 $10^{-4}$ 0.5 0.9 27.6 7.6 4.7 7.7 45.7 15.7
$\bm{10^{-4}}$ 1.0 11.7 45.5 55.2 4.2 18.0 2.7 22.9 $10^{-4}$ 1.0 1.4 27.9 8.2 5.9 8.9 46.5 16.5
XLS-R $10^{-3}$ 0.1 13.7 47.9 57.1 3.8 13.7 2.4 23.1 XLS-R $10^{-3}$ 0.1 0.4 29.0 8.2 4.3 7.9 50.5 16.7
$10^{-3}$ 0.5 13.0 50.7 56.2 3.8 14.6 2.7 23.5 $10^{-3}$ 0.5 0.9 29.1 11.1 4.7 8.4 52.1 17.7
$10^{-3}$ 1.0 12.6 45.4 58.0 3.9 14.8 2.6 22.9 $10^{-3}$ 1.0 1.4 27.7 12.0 4.3 8.9 49.4 17.3
$\bm{10^{-4}}$ 0.1 12.2 50.7 55.8 3.6 14.5 2.4 23.2 $10^{-4}$ 0.1 0.4 27.6 8.8 3.8 8.0 94.1 23.8
$10^{-4}$ 0.5 12.2 50.3 58.9 3.7 14.9 2.5 23.8 $\bm{10^{-4}}$ 0.5 0.6 31.6 9.8 4.3 8.3 45.2 16.6
$10^{-4}$ 1.0 13.5 49.2 58.0 3.7 15.2 2.8 23.7 $10^{-4}$ 1.0 0.8 31.2 10.7 4.5 9.5 47.1 17.3
Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ khm kmr kor nbl nno tat Avg Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ mrj slv snd sot spa urd Avg
# ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) # ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
3 MMS $10^{-3}$ 0.1 38.7 13.4 27.3 7.3 1.4 17.5 17.5 4 MMS $10^{-3}$ 0.1 9.3 10.3 19.5 12.5 4.9 35.4 15.3
$10^{-3}$ 0.5 37.5 14.9 27.1 8.0 2.2 19.3 18.2 $\bm{10^{-3}}$ 0.5 9.2 11.5 18.8 13.4 4.6 23.6 13.5
$10^{-3}$ 1.0 34.9 13.0 24.9 7.9 0.6 19.2 16.7 $10^{-3}$ 1.0 9.4 12.9 20.3 14.7 5.4 30.4 15.5
$\bm{10^{-4}}$ 0.1 34.3 12.6 25.5 7.4 0.8 16.8 16.2 $10^{-4}$ 0.1 8.4 9.3 21.4 11.8 4.7 27.0 13.8
$10^{-4}$ 0.5 35.3 13.8 26.6 8.3 0.8 20.8 17.6 $10^{-4}$ 0.5 8.9 10.2 19.0 13.1 4.4 33.6 14.9
$10^{-4}$ 1.0 36.8 13.5 26.4 7.9 0.5 20.1 17.5 $10^{-4}$ 1.0 9.6 12.5 18.7 13.6 4.4 27.4 14.4
XLS-R $10^{-3}$ 0.1 34.5 15.2 25.8 8.5 0.7 17.2 17.0 XLS-R $\bm{10^{-3}}$ 0.1 11.8 8.7 21.8 14.8 5.4 28.6 15.2
$10^{-3}$ 0.5 47.0 17.7 29.3 10.7 3.0 19.8 21.2 $10^{-3}$ 0.5 11.8 13.0 21.0 16.1 4.8 39.0 17.6
$10^{-3}$ 1.0 40.6 18.2 27.4 10.0 1.1 19.9 19.5 $10^{-3}$ 1.0 13.9 18.1 22.7 17.2 4.5 39.6 19.3
$\bm{10^{-4}}$ 0.1 33.1 14.9 29.9 9.3 2.8 19.5 18.2 $10^{-4}$ 0.1 12.3 9.3 21.9 14.2 4.6 34.9 16.2
$10^{-4}$ 0.5 43.6 16.4 27.8 9.4 1.1 22.7 20.2 $10^{-4}$ 0.5 14.5 13.9 23.7 17.5 5.5 40.7 19.3
$10^{-4}$ 1.0 46.0 19.6 28.3 10.7 2.3 23.5 21.7 $10^{-4}$ 1.0 12.8 13.2 20.8 15.0 4.4 30.4 16.1
Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ deu eng heb jpn rus spa Avg
# ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
5 MMS $10^{-3}$ 0.1 8.3 13.4 43.5 54.7 13.3 8.0 23.5
$10^{-3}$ 0.5 10.0 14.1 31.9 53.0 13.9 8.7 21.9
$\bm{10^{-3}}$ 1.0 12.2 15.6 41.9 52.4 14.3 9.8 24.4
$10^{-4}$ 0.1 8.2 14.8 32.9 64.2 14.1 8.7 23.8
$10^{-4}$ 0.5 9.8 15.3 39.0 65.6 14.6 9.4 25.6
$10^{-4}$ 1.0 12.6 16.8 38.0 74.5 14.9 12.8 28.3
XLS-R $10^{-3}$ 0.1 7.7 13.0 40.6 111.5 12.4 7.7 32.1
$10^{-3}$ 0.5 9.2 13.8 48.8 119.3 12.9 28.1 38.7
$10^{-3}$ 1.0 11.1 15.5 48.9 127.7 16.1 18.2 39.6
$\bm{10^{-4}}$ 0.1 6.1 11.2 41.5 77.1 11.1 8.9 26.0
$10^{-4}$ 0.5 9.6 13.0 45.4 105.5 11.9 8.3 32.3
$^{-4} 1.0 10$.9 14.1 44.9 118.8 12.3 9.0 35.0
Table 8: Results of the CTC-DRO models on the development set for the first two language sets using additional amounts of training data per language, where languages are indicated by their ISO code. We show the CER on the individual languages and CER averaged across languages (Avg) for fine-tuned MMS and XLS-R models. We highlight the best hyperparameter setting per set.
Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ ces cmn nan pol ron spa Avg Set Model $\bm{\eta_{q}}$ $\bm{\alpha}$ eng fas hrv ita slk yue Avg
# ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) # ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
1 MMS $10^{-3}$ 0.1 8.4 57.6 68.6 6.9 9.5 5.3 26.1 2 MMS $10^{-3}$ 0.1 9.4 20.5 8.1 7.2 10.8 53.4 18.2
$10^{-3}$ 0.5 8.0 48.6 64.8 7.0 9.6 5.4 23.9 $10^{-3}$ 0.5 9.6 20.4 8.8 7.5 11.3 52.4 18.3
$10^{-3}$ 1.0 8.5 50.8 71.5 7.5 9.8 5.2 25.5 $10^{-3}$ 1.0 9.5 19.5 8.9 7.5 10.8 49.8 17.6
$10^{-4}$ 0.1 8.1 50.1 64.0 6.6 9.7 5.2 24.0 $10^{-4}$ 0.1 9.6 18.8 8.6 7.5 10.5 55.1 18.4
$\bm{10^{-4}}$ 0.5 7.9 45.6 60.3 6.8 9.8 5.2 22.6 $10^{-4}$ 0.5 9.4 20.3 8.4 7.5 10.9 48.2 17.5
$10^{-4}$ 1.0 8.0 49.1 68.5 7.0 9.5 5.3 24.6 $\bm{10^{-4}}$ 1.0 9.4 19.9 8.9 7.4 11.3 47.8 17.5
XLS-R $10^{-3}$ 0.1 9.1 57.8 67.5 8.1 11.2 6.6 26.7 XLS-R $10^{-3}$ 0.1 11.6 24.6 10.2 9.0 13.4 56.9 21.0
$10^{-3}$ 0.5 12.9 57.8 69.8 10.4 13.2 7.8 28.7 $10^{-3}$ 0.5 11.7 22.7 9.7 8.2 12.9 57.9 20.5
$10^{-3}$ 1.0 11.1 53.3 67.2 9.3 12.7 7.5 26.9 $10^{-3}$ 1.0 23.2 30.7 18.4 15.3 21.7 83.0 32.1
$^{-4} 0.1 10$.6 61.4 70.1 9.3 11.6 6.9 28.3 $10^{-4}$ 0.1 11.5 25.7 10.1 8.0 12.8 91.0 26.5
$10^{-4}$ 0.5 12.7 56.7 69.7 10.0 13.5 8.0 28.4 $10^{-4}$ 0.5 19.2 27.0 16.3 12.6 18.9 68.6 27.1
$\bm{10^{-4}}$ 1.0 12.3 52.9 67.2 10.3 13.7 8.3 27.5 $\bm{10^{-4}}$ 1.0 11.6 25.1 9.6 9.1 14.4 50.3 20.0
F.2 Language-Specific Test Results
For each language set, we present the language-specific test set results of our experiments using balanced training data in Table 9. Table 10 shows the language-specific test set results for the first two sets based on experiments using all available training data in ML-SUPERB 2.0. In Table 11, we present results using WER on set 4 (balanced setup for brevity), which contains languages with clear word boundaries. Using this evaluation metric, CTC-DRO still achieves substantial worst-language improvements, namely 22.3% (MMS) and 11.8% (XLS-R) relative WER reductions. For MMS, the average WER is substantially reduced (14.4% relative). For XLS-R, the average WER increased marginally (0.4% relative), even though the average CER improved. This shows that character-level and word-level improvements do not always align, as a single character error invalidates an entire word. This also causes different languages to emerge as worst-performing under the CER versus the WER metrics. Despite the slight average WER increase for one model, CTC-DRO achieves its primary objective of substantially improving the performance on the worst-performing language.
Table 9: Results of the baseline models (Base), group DRO models (GDRO), and CTC-DRO models (Ours) on the test set for the different language sets, where languages are indicated by their ISO code. We show the CER on the individual languages, CER averaged across languages (Avg CER), and LID accuracy (LID) for fine-tuned MMS and XLS-R models. Best LID and CER results are highlighted, and the CERs for the worst-performing languages are underlined.
@@ F.2 Language-Specific Test Results
1 MMS Base 8.4 52.4 60.8 3.6 13.3 1.8 23.4 97.4
GDRO 20.6 48.6 86.6 4.3 16.7 6.2 30.5 78.7
Ours 10.5 46.1 56.8 3.7 17.9 2.3 22.9 95.8
XLS-R Base 7.3 64.9 60.8 3.1 13.4 1.8 25.2 92.6
GDRO 27.4 48.9 78.4 3.7 14.9 6.6 30.0 87.8
Ours 7.8 50.7 57.6 3.0 14.2 1.8 22.5 89.5
@@
2 MMS Base 0.2 21.8 9.0 5.9 8.2 49.4 15.8 98.4
GDRO 11.8 29.7 10.8 6.2 10.2 55.5 20.7 98.2
Ours 0.5 22.1 8.8 5.5 8.6 44.4 15.0 96.2
XLS-R Base 0.1 20.6 10.9 4.6 8.9 68.8 19.0 94.2
GDRO 12.7 28.5 14.4 5.1 10.2 58.8 21.6 87.0
Ours 0.5 21.5 12.6 5.2 10.0 45.0 15.8 89.3
@@
3 MMS Base 31.3 12.2 34.2 7.4 2.5 9.0 16.1 98.5
GDRO 33.2 19.1 34.0 22.4 9.8 13.5 22.0 98.7
Ours 31.3 12.0 27.6 8.1 2.3 10.2 15.3 98.7
XLS-R Base 33.2 13.3 32.3 8.7 3.7 11.0 17.0 99.2
GDRO 38.0 23.9 35.5 26.6 11.9 14.9 25.1 97.2
Ours 32.2 14.8 31.9 10.1 5.0 12.0 17.7 97.9
@@
4 MMS Base 14.8 6.9 24.0 14.4 5.9 20.1 14.4 87.9
GDRO 13.1 14.4 19.0 17.1 3.8 21.8 14.9 91.9
Ours 17.7 8.1 17.5 11.4 4.4 18.4 12.9 87.3
XLS-R Base 14.0 4.8 23.3 11.6 4.2 29.7 14.6 88.4
GDRO 19.5 25.6 18.5 23.0 3.9 21.1 18.6 83.5
Ours 11.9 6.7 21.0 13.8 4.8 24.2 13.7 88.9
@@
5 MMS Base 5.4 11.1 30.2 90.0 12.0 7.2 26.0 96.3
GDRO 27.6 27.0 32.6 62.2 17.6 8.4 29.2 67.0
Ours 10.9 15.4 39.2 57.5 13.2 9.3 24.3 90.5
XLS-R Base 4.8 9.2 33.2 114.8 10.5 7.1 29.9 89.0
GDRO 29.1 26.8 46.1 92.9 16.5 9.3 36.8 57.7
Ours 5.7 9.6 38.6 71.5 10.1 7.3 23.8 91.0
Table 10: Results of the baseline models (Base), group DRO models (GDRO), and CTC-DRO models (Ours) on the test set for the first two language sets using additional amounts of training data per language, where languages are indicated by their ISO code. We show the CER on the individual languages, CER averaged across languages (Avg CER), and LID accuracy (LID) for fine-tuned MMS and XLS-R models. Best LID and CER results are highlighted, and the CERs for the worst-performing languages are underlined.
@@
1 MMS Base 9.1 58.9 67.5 6.0 7.1 5.0 25.6 98.1
GDRO 13.8 92.1 96.3 6.7 11.9 5.8 37.8 83.9
Ours 8.7 45.9 62.8 6.2 7.5 5.3 22.8 98.5
XLS-R Base 13.0 92.1 78.3 9.8 12.0 8.5 35.6 96.4
GDRO 18.9 86.4 90.8 5.7 21.6 5.0 38.1 72.3
Ours 12.9 52.5 67.5 9.0 11.9 7.8 26.9 97.1
@@
2 MMS Base 9.6 16.9 8.5 6.8 8.0 66.9 19.5 99.0
GDRO 10.1 70.0 24.3 7.9 14.8 105.4 38.8 81.0
Ours 9.7 18.1 8.3 6.6 7.3 48.1 16.4 99.1
XLS-R Base 11.9 32.2 9.6 8.1 9.2 97.2 28.0 98.2
GDRO 8.8 88.2 33.9 6.7 23.3 102.9 44.0 80.8
Ours 11.6 23.2 9.3 8.2 8.9 51.4 18.8 98.6
Table 11: Results of the baseline models (Base), group DRO models (GDRO), and CTC-DRO models (Ours) on the test set for set 4, where languages are indicated by their ISO code. We show the WER on the individual languages, WER averaged across languages (Avg WER), and LID accuracy (LID) for fine-tuned MMS and XLS-R models. Best LID and WER results are highlighted, and the WERs for the worst-performing languages are underlined.
Set # Model Type mrj slv snd sot spa urd Avg WER LID
4 MMS Base 59.2 32.4 65.9 52.1 30.1 56.4 49.4 87.9
GDRO 57.3 56.1 50.3 61.5 19.1 56.6 50.2 91.9
Ours 51.2 36.7 49.4 43.6 22.5 50.3 42.3 87.3
XLS-R Base 60.2 22.9 63.9 44.6 21.4 74.0 47.8 88.4
GDRO 71.3 82.5 51.0 75.8 19.8 57.2 59.6 83.5
Ours 58.8 29.2 59.5 51.0 24.1 65.3 48.0 88.9
F.3 Ablation Study